OpenStudy (anonymous):

Find a vector of magnitude 5 units, and parallel to the resultant of the vectors a = 2i + 3j + k and b = i - 2j + k

OpenStudy (anonymous):

is the answer 5( a/root a^2 + b^2 + c^2)?

ganeshie8 (ganeshie8):

find the unit vector along a+b, then multiply it by 5

OpenStudy (anonymous):

parallel vectors have x1/x2 = y1/y2 = z1/z2

ganeshie8 (ganeshie8):

a = 2i + 3j + k b = i - 2j + k -------------------- a+b = 3i + j + 2k

ganeshie8 (ganeshie8):

thats the resultant of a and b, right ?

OpenStudy (anonymous):

oh okay so first I find a+b's unit vector, multiply with 5.. that's the magnitude.. To find if the vectors are parallel, I put its direction ratios in the form of x1/x2 = y1/y2 = z1/z2 if its equal, its parallel.. Am I right?

ganeshie8 (ganeshie8):

thats right, but we can look at it another way : since you're just multiplying by 5, which is a scalar, the vector will be parallel to the unit vector anyways.

OpenStudy (anonymous):

soo unit vector is..  3/root14, 1/root14, 2/root14 multiplied by 5 = 15/root14, 5/roo14, 10/root14